Please help. I'm having trouble with this problem.
Suppose you travel north for 65 kilometers then travel east 75 kilometers. How far are you from your starting point?
Answer:
Show work in this space.




Answer by venugopalramana(3286)

Please help. I'm having trouble with this problem.
Suppose you travel north for 65 kilometers then travel east 75 kilometers. How far are you from your starting point?
Answer:
Show work in this space.
ANGLE BETWEEN THE 2 DIRECTIONS OF TRAVEL =90
HENCE USING PYTHOGARUS THEOREM...
65^2+75^2=HYPOTENUSE^2
=4225+5625=H^2
=9850=H^2
H=99.2
HE IS AT 99.2 KM FROM STARTING POINT